# OmniSVG: A Unified Scalable Vector Graphics Generation Model

### v2

@@ Abstract
Scalable Vector Graphics (SVG) is an important image format widely adopted in graphic design because of their resolution independence and editability. The study of generating high-quality SVG has continuously drawn attention from both designers and researchers in the AIGC community. However, existing methods either produces unstructured outputs with huge computational cost or is limited to generating monochrome icons of over-simplified structures. To produce high-quality and complex SVG, we propose OmniSVG, a unified framework that leverages pre-trained Vision-Language Models (VLMs) for end-to-end multimodal SVG generation. By parameterizing SVG commands and coordinates into discrete tokens, OmniSVG decouples structural logic from low-level geometry for efficient training while maintaining the expressiveness of complex SVG structure. To further advance the development of SVG synthesis, we introduce MMSVG-2M, a multimodal dataset with two million richly annotated SVG assets, along with a standardized evaluation protocol for conditional SVG generation tasks. Extensive experiments show that OmniSVG outperforms existing methods and demonstrates its potential for integration into professional SVG design workflows.
Figure 1: Highlighted features of OmniSVG. OmniSVG is capable of autoregressively generating high-quality SVGs across a wide spectrum of complexity — from simple icons to intricate anime characters. It demonstrates remarkable versatility through multiple generation modalities, including Text-to-SVG, Image-to-SVG, and Character-Reference SVG, making it a powerful and flexible solution for diverse creative tasks.
@@ 1 Introduction
Existing methods adopt either optimization-based methods or auto-regressive approaches to generate SVG contents. The optimization-based methods [21, 33, 59] iteratively refine the SVG parameters by optimizing differentiable vector graphics rasterizers. Though these methods are effective in generating SVG icons, they suffer from significant computational overhead when scaling up to intricate samples and produce unstructured outputs with redundant anchor points. In contrast, auto-regressive methods build transformer models or adapt pre-trained Large Language Models (LLMs) to directly generate XML parameters [58] or codes [56, 41] representing SVGs. Benefiting from the end-to-end learning pipeline, the auto-regressive method is a more scalable approach [5] as it can learn directly from a large collection of SVG samples. However, existing auto-regressive approaches are limited to basic SVG contents [11, 22, 52] because of the limited context length and the scarcity of complex SVG data.
In this paper, we propose OmniSVG that harnesses native VLMs [1] for various end-to-end multimodal SVG generation tasks. By parameterizing SVG coordinates and commands into discrete tokens, OmniSVG decouples structural logic from low-level geometry, mitigating the “coordinate hallucination” problem prevalent in code-based LLMs, and produces vivid and colorful SVG results. Additionally, the next token prediction training objective enables OmniSVG to complete SVGs with diverse generation results given some partial observations. Compared to traditional auto-regressive SVG generation methods, OmniSVG is able to parameterize SVGs exceeding $30k$ tokens, facilitating the generation of detailed and complex SVG contents. Building upon pre-trained VLMs, our method natively integrates the ability to reason upon visual and textual instructions to synthesize editable, high-fidelity SVGs across diverse domains, from icons to intricate illustrations and anime characters.
To advance the development of SVG synthesis, we introduce MMSVG-2M, a multi-modal SVG synthesis dataset with two million richly annotated assets, encompassing icons, illustrations, and anime designs. We also establish a standardized evaluation protocol, MMSVG-Bench, for three key SVG generation tasks, i.e. “Text-to-SVG”, “Image-to-SVG”, and “Character Reference SVG Generation”. Extensive experiments show that OmniSVG produces highly detailed and complex SVG contents, surpassing prior art both quantitatively and qualitatively.
@@ 2 Related Works
SVG Generation. Early attempts to generating SVGs directly utilize architectures like RNNs [17, 40, 18, 43, 44], VAEs [4, 31, 48, 45, 50, 65], and Transformers [4, 55] to compress SVG commands into latent representations. Meanwhile, DeepSVG [4] further parameterizes SVGs using a dual transformer architecture but struggles with geometric consistency. Recently, the advent of large language models (LLMs) [29, 63, 51, 60, 5, 6, 62, 61, 47] unleashes the potential of generating SVGs via XML code synthesis [58, 56, 41]. However, the limited context length of existing LLM-based SVG generation methods [56, 41, 58] poses significant challenges in handling complex SVGs that exceed $10k$ tokens. In this paper, we explore the potential of native Vision-Language Models (VLMs) in multi-modal SVG generation. Combining pre-trained VLMs with SVG command parameterization, we validate that OmniSVG is able to follow multi-modal instructions and generate more complex SVGs than before.
Image Vectorization. Recent advancements in vectorization harness diffusion models paired with differentiable rasterizers, using techniques like score distillation sampling [36, 21, 7] and specialized regularizers [27, 33] to convert raster images into SVG paths. While these methods achieve remarkable results, they face limitations such as over-smoothing, color over-saturation, and lack of editability, often producing tangled paths that fail to capture hierarchical structures inherent in professional SVG designs. In this paper, we present an end-to-end generation method that directly follows multi-modal instructions to generate high-quality SVGs with improved path clarity.
SVG Datasets and Benchmarks. The lack of suitable datasets for complex SVG structures presents a significant challenge. Existing datasets [11, 22, 52] primarily focus on simplified path-based SVGs or monochrome icons, overlooking the intricate layered structures and rich color semantics found in real-world designs. For example, FIGR-8-SVG [11] focuses on monochromatic icons, while StarVector [41] proposes categorized datasets, including illustrations, icons, emojis, and fonts. Therefore, existing datasets only present simple SVG samples that do not exceed $8.2k$ tokens, failing to capture the complexities of layered structures and rich color semantics. Benchmark evaluations, such as VGBench [71], further highlight gaps in multi-format testing and the absence of comprehensive coverage for illustrative SVGs. To this end, we introduce MMSVG-2M, a multimodal SVG synthesis dataset comprising two million richly annotated assets, including icons, illustrations, and complex anime designs. We also present a standardized evaluation protocol, MMSVG-Bench, to evaluate diverse multi-modal SVG generation tasks with varying complexity.
@@ 3 OmniSVG Dataset
MMSVG-2M is a large-scale SVG dataset with two million SVG samples covering website icons, illustrations, graphic designs, anime characters, and so on (Sec. 3.1). To further promote the downstream development of SVG generation methods, we also introduce MMSVG-Bench, a benchmark consisting of a series of multi-modal instruction following tasks for conditional SVG generation (Sec. 3.2).
@@ 3.1 MMSVG-2M
The raw SVG we collected from the Internet consists of three types: MMSVG-Icon sourced from Iconfont, SVG illustrations sourced from iconsount, and SVG anime characters sourced from Freepik and our data creation pipeline depicted in Fig. 2. All these websites are online platforms where users can publish and share SVGs, encompassing a broad variety of categories. Specifically, our dataset comprises 2 million SVG samples within various subsets, including $1.1$ million SVG icons, $0.5$ million SVG illustrations, and 0.4 million SVG anime characters. We provide the data statistics of MMSVG-2M in Tab. 1.
Dataset Train Val Test Source Token Length
MMSVG-Icon 990k 106.7k 3.3k Iconfonthttps://www.iconfont.cn/ 2.2k $\pm$ 0.9k
MMSVG-Illustration 450k 48.5k 1.5k IconSount 8.1k $\pm$ 3.3k
MMSVG-Character 350k 48.9k 1.1k Freepik & generated 28k $\pm$ 7.3k
Table 2: SVG draw-commands. Draw commands used in this work along with their arguments and a visualization are listed. The start-position $(x_{1}$, $y_{1})$ is implicitly defined as the end-position of the preceding command.
@@
$x_{2}$, $y_{2}$ Move the cursor to the end-point $(x_{2},y_{2})$ without drawing anything. [Uncaptioned image]
@@
F
(Fill) $fill$ Draw the fill attribute of the path. $\varnothing$
@@
Data Curation.
We extract SVG samples with a comprehensive deduplication process based on filenames, SVG code, and metadata. We first fit the collected SVGs within a viewbox of $200\times 200$. Then, we employ an off-the-shelf VLM, specifically BLIP-2 [26], to generate captions for the SVGs. Please find more samples from the MMSVG-2M dataset in Fig. 10, and instruction templates in Sec. A.2.
SVG Simplification
is an essential procedure in SVG data cleansing, since the over-complicated XML grammars in the crawled SVG data will lead to ambiguities while representing basic shapes. For example, the Rect command creates a rectangular shape controlled by the starting point, width, and height arguments, like <rect x="80" y="90" width="100" height="100"/>, which can also be represented with four orthogonal lines. Also, the Circle Command creates a circle shape based on a center point and a radius, like <circle cx="50" cy="50" r="50"/>, which can also be approximated with the Bézier curve. Moreover, the Transform attribute in XML grammars applies an affine transformation to an existing shape. For example, <rect x="50" y="5" width="40" height="40" fill="blue" transform="rotate(45)"/> refers to rotating the blue rectangle with $45$ degrees.
To standardize training and evaluation, we simplify all SVG commands with atomic commands as shown in Tab. 2. Inspired by FIGR-8-SVG [11] and IconShop [55], we remove all attributes and simplify each SVG with five basic commands, including “Move To” (M), “Line To” (L), “Cubic Bézier” (C), “Elliptical Arc” (A), “ClosePath” (Z). The introduction of atomic commands further removes the ambiguities, as complex XML grammars can be approximated with the combination of several atomic commands. To efficiently produce a unified and less complex data structure, we utilize picosvg https://github.com/googlefonts/picosvg to remove grammars like “group” and “transform”, and simplify the complex commands to atomic path commands. It is worth noting that atomic path commands are sufficient to represent complex SVGs shown in Fig. 1.
@@ 3.2 MMSVG-Bench
As a multi-modal SVG generation evaluation protocol, MMSVG-Bench put emphasis on both the generation quality and how well the model follows multi-modal instructions.
Text-to-SVG requires a model to generate SVGs from text instructions. We measure the visual quality with Frechet Inception Distance (FID) [49], aesthetic appeal with Aesthetic score [42], text-SVG alignment with CLIP score [38], and Human Preference Scores (HPS) [57].
Image-to-SVG evaluates a model’s ability to convert images into SVGs. To quantify the distance between the input and output SVG, we calculate the cosine similarity of DinoV2 features (DinoScore) [34], Structural Similarity Index (SSIM) [53], Learned Perceptual Image Patch Similarity (LPIPS) [66], and Mean Squared Error (MSE).
Character-Reference SVG Generation evaluates a model’s ability to generate novel SVGs while keeping the profile of the characters depicted in the input image. Different from image-to-SVG, the model does not reconstruct, but generates a specific character SVG for the input image (see Fig. 6).
We evaluate the alignment between input character images and generated SVGs by prompting GPT-4o [20] to generate a score ranging from 1 to 10, the higher the better. For evaluation metrics, we also calculate the average token length (# tokens) of a generated SVG sample utilizing the Qwen2.5-VL [1] tokenizer. For computational cost, we compute the average generation time of generating per SVG sample.
@@ 4 OmniSVG
To support end-to-end training for multi-modal SVG generation, OmniSVG parameterizes a series of atomic path commands representing an SVG into a sequence before feeding into a pre-trained VLM with multi-modal instructions.
@@
SVG Tokenizer. As illustrated in Sec. 3, our MMSVG-2M dataset simplifies an SVG by removing all attributes and using five basic path commands (see Tab. 2). After the simplification, an SVG script $G$ is represented as the combination of $M$ paths, $G=\{P_{i}\}_{i=1}^{M}$. Here, $P_{i}$ is the $i$-th path containing $N_{i}$ commands, $P_{i}=\{C_{i}^{j}\}_{j=1}^{N_{i}}$, where $C_{i}^{j}$ is the $j$-th command in the $i$-th path. Each command is represented as $C_{i}^{j}=(U_{i}^{j},V_{i}^{j})$, containing both the type identifier $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z}\}$ and the corresponding location argument $V_{i}^{j}$. To generate colored SVG contents, we assign special tokens for hex values to control the “Fill” ($\mathrm{F}$) attribute, distinguishing it from the original SVG commands and coordinates. To this end, we are able to use a total six types of commands $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$ to parameterize a colored SVG parameterization. Specifically, our SVG tokenizer transforms SVG scripts $X_{s}$ into an ordered SVG token sequence within the same representation space as the pre-trained VLM. Following IconShop [55], we flatten the layered structure of the SVG script by concatenating different paths into a single command sequence, where each path begins with the drawing commands followed by point coordinates. Therefore, each SVG sequence could be represented as a flattened sequence. As the generation identifier, we apply special tokens like <SOP> and <EOS> to the two ends of a SVG sequence, identifying the begining and ending of a SVG sequence. We assign special tokens for each command type, i.e. $\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$. To shorten the length of the SVG sequence, we further merge the point coordinates into one token with a mapping function: $<x,y>\rightarrow x\times w+y$, where $w$ is the width of the image. The SVG sequence are then lifted into the same embedding space as the pre-trained VLM with a learnable embedding layer.
Model Architecture. OmniSVG leverages the Qwen2.5-VL architecture, a vision-language models (VLMs), which excels in processing both visual and textual inputs, enabling the generation of precise and compact SVG outputs. OmniSVG is trained for text-to-SVG, image-to-SVG, and character-reference SVG generation task. As depicted in Fig. 2, We first tokenize and embed the interleaved text and image inputs as prefix tokens. Then, we tokenize the SVG scripts into a sequence and concatenate it to the end of the prefix token with the SVG tokenizer. The full sequence is then served as the input of the decoder-only language model.
Training Objective. Similar to LLMs, we train the model to generate new tokens conditioned on the provided prefix tokens with the next-token prediction loss.
$\theta^{*}=\arg\max_{\theta}\prod_{i=1}^{L}P\left(x_{s,i}\mid x_{s,<i},x_{c}\right)$ (1)
Scaling Up. To study the effectiveness of scaling up multimodal SVG generation, we scale up OmniSVG from 3B to 7B parameters. We present training perplexity in Fig. 3, where both models are trained from scratch on 250 billion tokens. We show that, as the size of the model grows, the model achieves a lower validation perplexity, indicating a higher probability of producing the validation data.
(a) Training PPL for our models.
(b) Validation PPL for our models.

### v3

@@ Abstract
Scalable Vector Graphics (SVG) is an important image format widely adopted in graphic design because of their resolution independence and editability. The development of autonomous SVG generation workflows is continuously drawing attention from both designers and researchers in the AIGC community. However, existing methods either produce unstructured outputs at huge computational cost or are limited to generating monochrome icons of over-simplified structures. To produce high-quality and complex SVG adhering to multi-modal instructions, we propose OmniSVG, a unified SVG generation framework that inherits knowledge from a pre-trained Vision-Language Model (VLM). By parameterizing SVG commands and coordinates into discrete token sequences, the auto-regressive nature enables us to seaminglessly adapt modern VLMs to the direct SVG generation. To further advance the development of SVG synthesis, we introduce MMSVG-2M, a multimodal dataset with two million richly annotated SVG assets, along with a standardized evaluation protocol for conditional SVG generation tasks. Extensive experiments show that OmniSVG outperforms existing methods and demonstrates its potential for integration into professional SVG design workflows.
@@
Figure 1: OmniSVG is capable of autoregressively generating high-quality Scalable Vector Graphs (SVG) across a wide spectrum of complexity, from simple icons to intricate anime characters. OmniSVG demonstrates remarkable versatility in generating high-quality SVGs adhering to multimodal instructions, covering tasks like Text-to-SVG, Image-to-SVG, and Character-Reference SVG, making it a powerful and flexible solution for diverse creative tasks.
@@ 1 Introduction
Existing methods adopt either optimization-based methods or auto-regressive approaches to generate SVG contents.
The optimization-based methods (live; vtracer; diffvg) iteratively refine the SVG parameters by minimizing the differences between the input image and the raster image created by differentiable vector graphics rasterizers. Though these methods are sufficient for reconstructing SVG icons, they suffer from significant computational overhead when scaling up to more intricate samples and produce unstructured outputs with redundant anchor points, harming the editability of the reconstructed SVG samples. In contrast, auto-regressive methods build transformer models or adapt pre-trained Large Language Models (LLMs) to directly generate XML parameters (llm4svg) or codes (chat2svg; starvector) representing SVGs. Benefiting from the end-to-end learning pipeline, the auto-regressive method is a more scalable approach (meshxl) as it can learn directly from a large collection of SVG samples. However, existing auto-regressive approaches are limited to basic SVG contents (figr; stack; deepvecfont) because of the limited context length and the scarcity of complex SVG data.
In this paper, we propose OmniSVG that harnesses native VLMs qwen2.5-VL for various end-to-end multimodal SVG generation tasks. By parameterizing SVG coordinates and commands into discrete tokens, OmniSVG decouples structural logic from low-level geometry, mitigating the “coordinate hallucination” problem prevalent in code-based LLMs, and produces vivid and colorful SVG results. Additionally, the next token prediction training objective enables OmniSVG to complete SVGs with diverse generation results given some partial observations. Compared to traditional auto-regressive SVG generation methods, OmniSVG is able to parameterize SVGs exceeding $30k$ tokens, facilitating the generation of detailed and complex SVG contents. Building upon pre-trained VLMs, our method natively integrates the ability to reason upon visual and textual instructions to synthesize editable, high-fidelity SVGs across diverse domains, from icons to intricate illustrations and anime characters.
To advance the development of SVG synthesis, we introduce MMSVG-2M, a multi-modal SVG synthesis dataset with two million richly annotated assets, encompassing icons, illustrations, and anime designs.
We also establish a standardized evaluation protocol, MMSVG-Bench, for “Text-to-SVG” and “Image-to-SVG” generation. Extensive experiments show that OmniSVG can produce highly detailed and complex SVG contents, surpassing prior art both quantitatively and qualitatively.
@@ 2 Related Works
SVG Generation. Early attempts to generating SVGs directly utilize architectures like RNNs (sketchrnn_david_2018; im2vec; supersvg_hu_2024; ClipGen_Shen_2022; CLIPVG_song_2023), VAEs (deepsvg; fonts_svg; strokenuwa_tang_2024; marvel_su_2023; evolution_tian_2022), and Transformers (deepsvg; iconshop) to compress SVG commands into latent representations. Meanwhile, DeepSVG (deepsvg) further parameterizes SVGs using a dual transformer architecture but struggles with geometric consistency. Recently, the advent of large language models (LLMs) (llava_Liu_2023; yi_young_2024; qwen2_wang_2024; qwen2.5; meshxl; ll3da; shapegpt; pminr; strokenuwa) unleashes the potential of generating SVGs via XML code synthesis (llm4svg; chat2svg; starvector). However, the limited context length of existing LLM-based SVG generation methods (chat2svg; starvector; llm4svg) poses significant challenges in handling complex SVGs that exceed $10k$ tokens. In this paper, we explore the potential of native Vision-Language Models (VLMs) in multi-modal SVG generation. By combining pre-trained VLMs with SVG command parameterization, we validate that OmniSVG is able to follow multi-modal instructions and generate vivid and complex SVGs.
Image Vectorization. Recent advancements in vectorization harness diffusion models paired with differentiable rasterizers, using techniques like score distillation sampling (dreamfusion; vectorfusion; svgbuilder) and specialized regularizers diffvg; live to convert raster images into SVG paths. While these methods achieve remarkable results, they face limitations such as over-smoothing, color over-saturation, and lack of editability, often producing tangled paths that fail to capture hierarchical structures inherent in professional SVG designs. In this paper, we present an end-to-end approach that follows multi-modal instructions to generate high-quality SVGs with improved path clarity and editability.
SVG Datasets and Benchmarks. The lack of suitable datasets for complex SVG structures presents a significant challenge. Existing datasets (figr; stack; deepvecfont) primarily focus on simplified path-based SVGs or monochrome icons, overlooking the intricate layered structures and rich color semantics found in real-world designs. For example, FIGR-8-SVG (figr) focuses on monochromatic icons, while StarVector (starvector) proposes categorized datasets, including illustrations, icons, emojis, and fonts. Therefore, existing datasets only present simple SVG samples that do not exceed $8.2k$ tokens, failing to capture the complexities of layered structures and rich color semantics. Benchmark evaluations, such as VGBench (vgbench), further highlight gaps in multi-format testing and the absence of comprehensive coverage for illustrative SVGs. To this end, we introduce MMSVG-2M, a multimodal SVG synthesis dataset comprising two million richly annotated assets, including icons, illustrations, and complex anime designs. We also present a standardized evaluation protocol, MMSVG-Bench, to evaluate diverse multi-modal SVG generation tasks with varying complexity.
@@ 3 OmniSVG Dataset
We present MMSVG-2M, a large-scale SVG dataset with two million SVG samples covering website icons, illustrations, graphic designs, anime characters, and etc (Sec.˜3.1). To promote the downstream development of SVG generation methods, we also introduce MMSVG-Bench, a standardized evaluation protocol for a series of multi-modal instruction following tasks for conditional SVG generation (Sec.˜3.2).
@@ 3.1 MMSVG-2M
With increasing visual complexity, MMSVG-2M consists of three subsets, 1) the icon subset MMSVG-Icon collected from Iconfont, 2) the illustration subset MMSVG-Illustration sourced from IconSount, and 3) the complex anime character subset MMSVG-Character both curated from Freepik and created by our data creation pipeline as shown in Fig.˜2. All these websites are online platforms where users can publish and share SVGs, encompassing a broad variety of categories. Specifically, our collection of MMSVG-2M contains $1.1$ million icons, $0.5$ million illustrations, and 0.4 million anime characters as shown in Tab.˜6.
Data Curation.
We extract SVG samples with a comprehensive deduplication process based on filenames, SVG code, and metadata. We first fit the collected SVGs within a viewbox of $200\times 200$. Then, we employ an off-the-shelf VLM, specifically BLIP-2 (blip2), to generate captions for the SVGs. Please find more samples from the MMSVG-2M dataset in Fig.˜8, and instruction templates in Sec.˜A.2.
SVG Simplification
is an essential procedure in SVG data cleansing, since the over-complicated XML grammars in the crawled SVG data will lead to ambiguities while representing basic shapes. To standardize training and evaluation, we simplify all SVG commands with atomic commands as shown in Tab.˜1. Inspired by FIGR-8-SVG figr and IconShop (iconshop), we remove all attributes and simplify each SVG with five basic commands, including “Move To” (M), “Line To” (L), “Cubic Bézier” (C), “Elliptical Arc” (A), “ClosePath” (Z). The introduction of atomic commands further removes the ambiguities, as complex XML grammars can be approximated with the combination of several atomic commands. To efficiently produce a unified and less complex data structure, we utilize picosvg https://github.com/googlefonts/picosvg to remove grammars like “group” and “transform”, and simplify the complex commands to atomic path commands. It is worth noting that atomic path commands are sufficient to represent complex SVGs shown in Fig.˜1.
Table 1: SVG Draw Commands. Draw commands used in this work along with their arguments and a visualization are listed. The start-position $(x_{1}$, $y_{1})$ is implicitly defined as the end-position of the preceding command.
@@
$x_{2}$, $y_{2}$
Move the cursor to the end-point $(x_{2},y_{2})$
without drawing anything.
@@
F (Fill) fillfill
Draw the fill attribute of the path.
$\varnothing$
@@ 3.2 MMSVG-Bench
To compensate for the vacancy of standardized and open evaluation for SVG generation, we introduce MMSVG-Bench, a comprehensive benchmark for multi-modal SVG generation. We require the corresponding benchmark to be a sufficient verification whether a model is practically useful in real-world scenarios, and avoid the excessive similarity between the benchmark input data and training data as in traditional train/test splits. Therefore, we opt to generate the benchmark inputs with GPT-4o. Specifically, for Text-to-SVG task, we synthesize 150 textual prompts for each SVG subset (i.e. Icon and Illustration). For Image-to-SVG task, we synthesize extra 150 textual descriptions, and prompt GPT-4o to generate vector-style images with transparent backgrounds based on the above texts as the ground truth visual samples. We focus on both the visual quality and semantics of the generation results.
Text-to-SVG
requires a model to generate SVGs from text instructions. We measure the visual quality with Frechet Inception Distance (FID) (fid), aesthetic appeal with Aesthetic score (aesthetic), text-SVG alignment with CLIP score (clip_score), and Human Preference Scores (HPS) (hps).
Image-to-SVG
evaluates a model’s ability to convert images into SVGs. To quantify the distance between the input and output SVG, we calculate the cosine similarity of DinoV2 features (DinoScore) (dinov2), Structural Similarity Index (SSIM) (ssim), Learned Perceptual Image Patch Similarity (LPIPS) (lpips), and Mean Squared Error (MSE).
Character-Reference SVG Generation
evaluates a model’s ability to generate novel SVGs while keeping the profile of the characters depicted in the input image. Different from image-to-SVG, the model does not reconstruct, but generates a specific character SVG for the input image (see Fig.˜5). We evaluate the alignment between input character images and generated SVGs by prompting GPT-4o (GPT-4o) to generate a score ranging from 1 to 10, the higher the better.  gloeckle2024better; katharopoulos2020transformers; guo2025log
@@ 4 OmniSVG
To support end-to-end training for multi-modal SVG generation, OmniSVG parameterizes a series of atomic SVG path commands into a sequence before feeding into a pre-trained VLM with multi-modal instructions.
SVG Tokenizer. As illustrated in Sec. 3, our MMSVG-2M dataset simplifies an SVG by removing all attributes and using five basic path commands (see Tab.˜1). After the simplification, an SVG script $G$ is represented as the combination of $M$ paths, $G=\{P_{i}\}_{i=1}^{M}$. Here, $P_{i}$ is the $i$-th path containing $N_{i}$ commands, $P_{i}=\{C_{i}^{j}\}_{j=1}^{N_{i}}$, where $C_{i}^{j}$ is the $j$-th command in the $i$-th path. Each command is represented as $C_{i}^{j}=(U_{i}^{j},V_{i}^{j})$, containing both the command type identifier $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z}\}$ and the corresponding location argument $V_{i}^{j}$.
To generate colored SVG contents, we assign special tokens for hex values to control the “Fill” ($\mathrm{F}$) attribute, distinguishing it from the original SVG commands and coordinates. To this end, we are able to use a total six types of commands $U_{i}^{j}\in\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$ to parameterize a colored SVG parameterization.
@@
Specifically, our SVG tokenizer transforms SVG scripts $X_{s}$ into an ordered SVG token sequence within the same representation space as the pre-trained VLM. Following IconShop (iconshop), we flatten the layered structure of the SVG script by concatenating different paths into a single command sequence, where each path begins with the drawing commands followed by point coordinates. Therefore, each SVG sequence could be represented as a flattened sequence. As the generation identifier, we apply special tokens like <SOP> and <EOS> to the two ends of a SVG sequence, identifying the begining and ending of a SVG sequence. We assign special tokens for each command type, i.e. $\{\mathrm{M},\mathrm{L},\mathrm{C},\mathrm{A},\mathrm{Z},\mathrm{F}\}$. To shorten the length of the SVG sequence, we further merge the 2D point coordinates into one token with a mapping function: $x,y>\rightarrow x\times w+y$, where $w$ is the width of the image. The SVG sequence are then lifted into the same embedding space as the pre-trained VLM with a learnable embedding layer.
Model Architecture. OmniSVG adopts Qwen2.5-VL qwen2.5-VL, an open-sourced VLM that excels at understanding intricate vision-text inputs, as its backbone (Fig.˜2) to produce precise and compact SVG outputs. OmniSVG is trained to predict the SVG suffix tokens ($x_{s}$) conditioned on the mutli-modal instruction prefix tokens ($x_{c}$) with the standard next-token prediction objective.
$\vskip-11.38109pt\theta^{*}=\arg\max_{\theta}\prod_{i=1}^{L}P\left(x_{s,i}\mid x_{s,<i},x_{c}\right)$ (1)